Given series:

"n"

"\ufeffSn = \u03a3 \ufeff" "\ufeff((k+1)^2 - k^2)"

"\ufeffk=1"


which of the following statements are true?


  1. it is a telescoping series
  2. it sums to n(n+1)
  3. it sums to n(n+2)
  4. it sums to n(n+1) - n
  5. it sums to the same value as:

"n"

"\u03a3" "(2k + 1)"

 "k = 1"


6) it sums to the same value as:


"n + 1"

"\u03a3" "(2k)"

"k = 2"



1
Expert's answer
2021-05-31T18:21:55-0400

1.false

A telescoping series is a series whose general term can be written as the difference of two consecutive terms of a sequence:

"t_n=a_n-a_{n+1}"


2.false

"S_n=\sum((k+1)^2-k^2)=(2^2-1)+(3^2-2^2)+(4^2-3^2)+...+((n+1)^2-n^2)="

"=(n+1)^2-1=n^2+2n=n(n+2)"


3.true


4.false


5.true

"S_n=\sum((k+1)^2-k^2)=\sum(k^2+2k+1-k^2)=\sum(2k+1)"


6.false

"\displaystyle{\sum_2^{n+1}2k}=4+8+10+...+2(n+1)=n(4+2n+2)/2=n(n+3)"